Assume that 𝛼 is increasing on [π‘Ž, 𝑏]. Note that if 𝑓 ∈ 𝑅(𝛼) on [π‘Ž, 𝑏], then 𝑓





1
Expert's answer
2022-02-03T16:32:24-0500

… then, "\displaystyle
|f|\in R(\alpha) \text{ on }[a,b]\text{ and we have the inequality;}"

"\displaystyle
\left|\int_a^bf(x)\ d\alpha(x)\right|\leq\int_a^b|f(x)|\ d\alpha(x)."

"\displaystyle
\color{red}\text{This is a standard theorem.}"